You want to save for your newborn child’s college education. You can invest in a tax free bond which pay 7% annual interest rate, compounded continuously. How much should you invest (principle) today so that you will have $100,000 in 18 years?
Answer #1

We use the formula:
A=P(e)^rn
where
A=future value
P=present value
r=rate of interest
n=time period.

e=2.71828

100,000=P*(2.71828)^(0.07*18)

P=100,000/(2.71828)^(0.07*18)

=100,000*0.283654267

=$28365.43(Approx)